OpenStudy (anonymous):

factor 9x^2-12x+36 like this ( )( )

OpenStudy (anonymous):

May be something like this \(9x{}^2-12x+36= \left(x-\frac{2}{3} \left(1-2 i \sqrt{2}\right) \right) \left( x- \frac{2}{3} \left(1+2 i \sqrt{2}\right)\right) \)

OpenStudy (king):

is that i cap?

OpenStudy (anonymous):

i cap ?!

OpenStudy (king):

idk what is i?

OpenStudy (anonymous):

wasnt the answer I was looking for but thanks..

OpenStudy (anonymous):

\( \large i=\sqrt{-1} \)

OpenStudy (king):

complex number?

OpenStudy (anonymous):

no its just that a square root can never be a negative number so its just an imaginary number.. I think,,

OpenStudy (king):

its a complex number...and this is possible